Determine the critical values for the confidence interval for the population variance from the given values. Round your answers to three decimal places.

n=18 and α=0.05.
Answer #1

Solution :

Given that,

n = 18

df = 18 - 1 = 17

Critical value are :

2L = 2/2,df = 30.191

2R = 21 - /2,df = 7.564